OpenStudy (anonymous):

The pH of a fruit juice is 2.5. Find the hydronium ion concentration, [H3O+], of the juice. Use the formula pH=-log[H3O+]

OpenStudy (kirbykirby):

\[2.5=-\log[\text{H}_3\text{O}^+] \] This log is in base 10, so use the property of logs: \(a=\log_b(x)\) means \(b^a=x\) If your log is base 10, then \(b=10\) and we don't write the subscript, so \(a=\log x\) means \(10^a=x\) If the above formula is confusing, just try imagining that you are substituting \(\text{H}_3\text{O}^+\) by "\(x\)" so, \(2.5=-\log x\) \(-2.5=\log x\) Can you figure out the rest? I.e. how to solve for x?